# RLMiner: Finding the Most Frequent k-sized Subgraph via Reinforcement Learning

### v1

@@ 1. Introduction
Finding the most frequent induced subgraph in a labeled undirected target graph is a crucial but challenging step in characterizing the complex structure property of graphs, it has significant implications across various domains, from bioinformatics [6; 2] and chemistry [5] to social network analysis [17; 22]. The task is defined as finding the most frequent induced subgraph of size $k$ given a undirected labeled target graph. As shown in Figure 1, subgraph (b) is the most frequent induced subgraph of size $3$ in target graph (a) with frequency of $4$.
@@
Extensive experiments on real-world datasets demonstrate that our proposed $RLMiner$ can solve the problem with high performance while the running time of exact enumeration is much longer (more than $2000\times$ on harder target graphs) than the running time of RLMiner. Our contribution can be summarized as follows:
@@ 3.1. Problem Setup
Note that, two induced subgraphs of $G^{T}$ isomorphic to $G^{S}$ are considered different only if they have at least one node they do not share. For instance, in Figure 1, there are $6$ isomorphism mappings of subgraph (c) in target graph (a):
@@
where $v^{T}_{i}$ denotes the node in target graph (a). However, these mappings all share the same node sets $\{v^{T}_{1},v^{T}_{2},v^{T}_{3}\}$, thus we count the frequency of subgraph (c) in target graph (a) as $1$. Obviously, this is more intuitive than using a frequency count of $6$. Following the same principle, subgraph (b) has a frequency count of $4$ in target graph (a).
@@ 3.2. Downward Closure Property Frequency Measure
Obviously, a frequency count of $4$ is more intuitive than a frequency count of $2$. Also, the MNI of subgraph (c) is $3$, which has a higher MNI than subgraph (b). Therefore, in our problem, we use the most straightforward frequency measurement without DCP restriction as defined in Definition 1.
Note that, in the above example, there are at most $2$ nodes with label “B” in target graph (a), thus the MNI of subgraph (b) is always bounded by $2$, such Downward Closure Property (DCP) allows traditional methods to prune a significant amount of search space. However, our proposed RLMiner does not rely on the frequency measurement with DCP to prune the search space, it can approximately find the most frequent subgraph with time complexity linear to $k$.
@@ 4.2.1. Task-state-aware GNN
where $OneHot$ is the one-hot encoding function, $\bm{W}_{a}$ and $\bm{W}_{b}$ are learnable parameters, we map the elements of $C_{t}$ to representations $\bm{h}_{C_{t},0}$ and $\bm{h}_{C_{t},1}$, $\odot$ denotes the element-wise product, and $\bm{h}_{C_{t}}$ is the task representation.
@@
Task-state-aware GNN To encode the local graph structure for large $k$, we need to stack our task-state-aware message-passing layers with a large depth. However, GNN with a large depth often suffer from over-smoothing [23] issue. Given a $N$-layer task-state-aware GNN, we employ skip connection [14] to connect the input of the initial GNN layer and the output of the last GNN layer to alleviate the effect of over-smoothing. Moreover, different subgraph sizes $k$ necessitate different sizes of local graph structures and, consequently, a different number of GNN layers. While this could be handled by a task-aware component within a task-state-aware message-passing framework, relying solely on the final GNN layer may introduce bias, particularly when the model has not been well-trained. To better filter out redundant information, we extract the local graph structure from the last GNN layer according to the task representation using the gating mechanism. The final task-state-aware node representation is computed using a gated skip connection:
@@ 5.1. Dataset
We conduct experiments over real-world datasets obtained from TUDatasets to evaluate RLMiner. For each dataset, we aim to find induced subgraphs of size $k\in\{5,6,7,8,9\}$, since the search space for subgraph size $k=3$ or $4$ is relatively small, and the exact enumeration failed to scale beyond $k>9$. For simplicity, only connected graphs with $500>|V^{T}|>10$ are used as target graphs, and $80\%$ of target graphs are used for training, the remaining target graphs are used for evaluation. The statistics of the target graph datasets used in our experiments are summarized in Table 1.
@@ 5.2. Baseline methods
Gaston was designed for finding all the frequent subgraphs across a graph dataset with frequency higher than a given frequency threshold, where a subgraph is counted as 0 or $1$ in a single target graph. Therefore, to enumerate all the induced subgraphs in a single target graph, we set frequency threshold to $1$.
T-FSM-serial: T-FSM [^3] https://github.com/lyuheng/T-FSM [37] is an exact enumeration method for finding frequent subgraphs in a single labeled graph with $MNI$ higher than a given threshold. We use T-FSM-serial to find the induced subgraphs with the highest $MNI$, then use VF3 to count the frequency of each founded induced subgraphs to find the most frequent one, this is considered as an approximate method.
However, the official implementation of T-FSM-serial does not return the $MNI$ of each founded subgraph, and it is difficult to directly determine a proper threshold. Thus, we initially set the threshold to the highest possible $MNI$ and run T-FSM-serial, then iteratively reduce the threshold and re-run T-FSM-serial until we find the induced subgraphs with the highest $MNI$ for all subgraph sizes $k\in K$.
Rand: We replace the Actor network of RLMiner by a random action sampler to sample $100$ induced subgraphs for a given target graph, then use VF3 to count the frequency of each sampled subgraphs to find the most frequent one, this serves as an approximate method. We repeat and average the results over $5$ seeds.
@@ 5.4. Result
For smaller subgraph size $k=5$ and $6$, the search space is relatively small, Rand can achieve high approximation ratio by simple sampling. However, the search space increases exponentially as $k$ increases, RLMiner has a much better performance than other baseline methods for larger $k$ across all datasets, even for the harder target graphs (ENZYMES, PROTEINS). For simpler target graphs (COX2, BZR, DHFR), RLMiner can almost perfectly find the most frequent induced subgraph for all $k$.
@@ Appendix B Implementation Details
For both the Actor network and the Critic network of RLMiner, the hidden dimension is set to $256$, we adopt $9$ layers of task-state-aware message passing layers, and the number of heads is set to $4$ for MHA and PMA layers.
For the training of RLMiner, we employ Prioritized Experience Replay (PER) [28] as our replay buffers, where each subgraph size $k$ has an individual replay buffer with size $10^{-6}$, the prioritization exponent of PER is set to $0.2$ and $\beta$ of PER is set to 0.6. Moreover, we set discount factor $\gamma$ to $0.99$, initial $\alpha$ to $1$ and target network soft update factor $\tau$ to $10^{-2}$. We adopt Adam optimizer with learning rate $2.5\times 10^{-4}$. Before training, we randomly sample states from training target graphs and store into replay buffers for $50$ epochs. Next, for COX2, BZR and DHFR datasets, we train RLMiner with $75$ epochs. For ENZYMES dataset, we train RLMiner with $150$ epochs since ENZYMES dataset is considered to be harder target graphs. For PROTEINs dataset, although it is also considered to be harder target graphs, we train RLMIner with $75$ epochs to limit the training time. During training, we perform $2$ gradient steps to update the parameters of RLMiner after each termination of episode, where we sample $8$ experiences from each replay buffer for each gradient step. In addition, although we only aim to find subgraphs with size $k=5$ to $k=9$, we still train RLMiner with subgraph size from $k=3$ up to $k=9$, this allows RLMiner to learn basic knowledge from simpler tasks instead of directly learning from harder tasks.
@@ C.2. Ablation Study
Task-aware gated skip connection To measure the importance of gated skip connection as described in Equation 10, we create a variant model (RLMiner skip w/o gate) that employs skip connection without gating mechanism, the average approximation ratio on test graphs are shown in Figure 7. As mentioned in Section 4, different subgraph size $k$ requires different size of local graph structures, directly using the final GNN layer may introduce bias if the model is not well-trained. Although it is possible to be filtered out by the task-aware message-passing, using the gated skip connection could filter out redundant information in a more direct way. Therefore, RLMiner with gated skip connection can achieve better performance for tasks that are harder to be well-trained (subgraph size of $8$ and $9$). State-aware message-passing To evaluate the effectiveness of the proposed state-aware message-passing, we create a variant model (RLMiner w/o state-mp), which concatenates $\bm{x}_{t}$ with the node label matrix of target graph and pass into task-aware GNN. As shown in Figure 8, both approaches exhibit similar performance on test target graphs for smaller values of $k$ , while RLMiner with state-aware message-passing has better performance for larger values of $k$. As explained in Section 4.2, concatenating $\bm{x}_{t}$ with the node label matrix may introduce bias when comparing the extracted local graph structures to determine ”frequency.” However, nodes in the available action space are always unselected, meaning they share the same state value in $\bm{x}_{t}$. Consequently, this concatenation does not significantly impact the ranking of each action’s probability or value. Also, for smaller values of $k$, only a few nodes are selected, which does not lead to much bias when comparing local graph structures, and the bias increases as $k$ becomes larger. Therefore, there is no significant performance gap between two approaches for smaller values of $k$, but RLMiner with state-aware message passing performed better for larger value of $k$.

### v2

@@ 1. Introduction
Finding the most frequent induced subgraph in a labeled undirected target graph is a crucial but challenging step in characterizing the complex structure property of graphs, it has significant implications across various domains, from bioinformatics [6; 2] and chemistry [5] to social network analysis [17; 22]. The task is defined as finding the most frequent induced subgraph of size $k$ given a undirected labeled target graph. As shown in Figure 1, subgraph (b) is the most frequent induced subgraph of size 33 in target graph (a) with frequency of 44.
@@
Extensive experiments on real-world datasets demonstrate that our proposed RLMinerRLMiner can solve the problem with high performance while the running time of exact enumeration is much longer (more than $2000\times$ on harder target graphs) than the running time of RLMiner. Our contribution can be summarized as follows:
@@ 3.1. Problem Setup
Note that, two induced subgraphs of $G^{T}$ isomorphic to $G^{S}$ are considered different only if they have at least one node they do not share. For instance, in Figure 1, there are 66 isomorphism mappings of subgraph (c) in target graph (a):
@@
where $v^{T}_{i}$ denotes the node in target graph (a). However, these mappings all share the same node sets $\{v^{T}_{1},v^{T}_{2},v^{T}_{3}\}$, thus we count the frequency of subgraph (c) in target graph (a) as 11. Obviously, this is more intuitive than using a frequency count of 66. Following the same principle, subgraph (b) has a frequency count of 44 in target graph (a).
@@ 3.2. Downward Closure Property Frequency Measure
Obviously, a frequency count of 44 is more intuitive than a frequency count of 22. Also, the MNI of subgraph (c) is 33, which has a higher MNI than subgraph (b). Therefore, in our problem, we use the most straightforward frequency measurement without DCP restriction as defined in Definition 1.
Note that, in the above example, there are at most 22 nodes with label “B” in target graph (a), thus the MNI of subgraph (b) is always bounded by 22, such Downward Closure Property (DCP) allows traditional methods to prune a significant amount of search space. However, our proposed RLMiner does not rely on the frequency measurement with DCP to prune the search space, it can approximately find the most frequent subgraph with time complexity linear to $k$.
@@ 4.2.1. Task-state-aware GNN
where OneHotOneHot is the one-hot encoding function, $\bm{W}_{a}$ and $\bm{W}_{b}$ are learnable parameters, we map the elements of $C_{t}$ to representations $\bm{h}_{C_{t},0}$ and $\bm{h}_{C_{t},1}$, $\odot$ denotes the element-wise product, and $\bm{h}_{C_{t}}$ is the task representation.
@@
Task-state-aware GNN To encode the local graph structure for large $k$, we need to stack our task-state-aware message-passing layers with a large depth. However, GNN with a large depth often suffer from over-smoothing [23] issue. Given a NN-layer task-state-aware GNN, we employ skip connection [14] to connect the input of the initial GNN layer and the output of the last GNN layer to alleviate the effect of over-smoothing. Moreover, different subgraph sizes $k$ necessitate different sizes of local graph structures and, consequently, a different number of GNN layers. While this could be handled by a task-aware component within a task-state-aware message-passing framework, relying solely on the final GNN layer may introduce bias, particularly when the model has not been well-trained. To better filter out redundant information, we extract the local graph structure from the last GNN layer according to the task representation using the gating mechanism. The final task-state-aware node representation is computed using a gated skip connection:
@@ 5.1. Dataset
We conduct experiments over real-world datasets obtained from TUDatasets to evaluate RLMiner. For each dataset, we aim to find induced subgraphs of size $k\in\{5,6,7,8,9\}$, since the search space for subgraph size $k=3$ or 44 is relatively small, and the exact enumeration failed to scale beyond $k>9$. For simplicity, only connected graphs with $500>|V^{T}|>10$ are used as target graphs, and $80\%$ of target graphs are used for training, the remaining target graphs are used for evaluation. The statistics of the target graph datasets used in our experiments are summarized in Table 1.
@@ 5.2. Baseline methods
Gaston was designed for finding all the frequent subgraphs across a graph dataset with frequency higher than a given frequency threshold, where a subgraph is counted as 0 or 11 in a single target graph. Therefore, to enumerate all the induced subgraphs in a single target graph, we set frequency threshold to 11.
T-FSM-serial: T-FSM [^3] https://github.com/lyuheng/T-FSM [37] is an exact enumeration method for finding frequent subgraphs in a single labeled graph with MNIMNI higher than a given threshold. We use T-FSM-serial to find the induced subgraphs with the highest MNIMNI, then use VF3 to count the frequency of each founded induced subgraphs to find the most frequent one, this is considered as an approximate method.
However, the official implementation of T-FSM-serial does not return the MNIMNI of each founded subgraph, and it is difficult to directly determine a proper threshold. Thus, we initially set the threshold to the highest possible MNIMNI and run T-FSM-serial, then iteratively reduce the threshold and re-run T-FSM-serial until we find the induced subgraphs with the highest MNIMNI for all subgraph sizes $k\in K$.
Rand: We replace the Actor network of RLMiner by a random action sampler to sample $100$ induced subgraphs for a given target graph, then use VF3 to count the frequency of each sampled subgraphs to find the most frequent one, this serves as an approximate method. We repeat and average the results over 55 seeds.
@@ 5.4. Result
For smaller subgraph size $k=5$ and 66, the search space is relatively small, Rand can achieve high approximation ratio by simple sampling. However, the search space increases exponentially as $k$ increases, RLMiner has a much better performance than other baseline methods for larger $k$ across all datasets, even for the harder target graphs (ENZYMES, PROTEINS). For simpler target graphs (COX2, BZR, DHFR), RLMiner can almost perfectly find the most frequent induced subgraph for all $k$.
@@ Appendix B Implementation Details
For both the Actor network and the Critic network of RLMiner, the hidden dimension is set to $256$, we adopt 99 layers of task-state-aware message passing layers, and the number of heads is set to 44 for MHA and PMA layers.
For the training of RLMiner, we employ Prioritized Experience Replay (PER) [28] as our replay buffers, where each subgraph size $k$ has an individual replay buffer with size $10^{-6}$, the prioritization exponent of PER is set to $0.2$ and $\beta$ of PER is set to 0.6. Moreover, we set discount factor $\gamma$ to $0.99$, initial $\alpha$ to 11 and target network soft update factor $\tau$ to $10^{-2}$. We adopt Adam optimizer with learning rate $2.5\times 10^{-4}$. Before training, we randomly sample states from training target graphs and store into replay buffers for $50$ epochs. Next, for COX2, BZR and DHFR datasets, we train RLMiner with $75$ epochs. For ENZYMES dataset, we train RLMiner with $150$ epochs since ENZYMES dataset is considered to be harder target graphs. For PROTEINs dataset, although it is also considered to be harder target graphs, we train RLMIner with $75$ epochs to limit the training time. During training, we perform 22 gradient steps to update the parameters of RLMiner after each termination of episode, where we sample 88 experiences from each replay buffer for each gradient step. In addition, although we only aim to find subgraphs with size $k=5$ to $k=9$, we still train RLMiner with subgraph size from $k=3$ up to $k=9$, this allows RLMiner to learn basic knowledge from simpler tasks instead of directly learning from harder tasks.
@@ C.2. Ablation Study
Task-aware gated skip connection To measure the importance of gated skip connection as described in Equation 10, we create a variant model (RLMiner skip w/o gate) that employs skip connection without gating mechanism, the average approximation ratio on test graphs are shown in Figure 7. As mentioned in Section 4, different subgraph size $k$ requires different size of local graph structures, directly using the final GNN layer may introduce bias if the model is not well-trained. Although it is possible to be filtered out by the task-aware message-passing, using the gated skip connection could filter out redundant information in a more direct way. Therefore, RLMiner with gated skip connection can achieve better performance for tasks that are harder to be well-trained (subgraph size of 88 and 99). State-aware message-passing To evaluate the effectiveness of the proposed state-aware message-passing, we create a variant model (RLMiner w/o state-mp), which concatenates $\bm{x}_{t}$ with the node label matrix of target graph and pass into task-aware GNN. As shown in Figure 8, both approaches exhibit similar performance on test target graphs for smaller values of $k$ , while RLMiner with state-aware message-passing has better performance for larger values of $k$. As explained in Section 4.2, concatenating $\bm{x}_{t}$ with the node label matrix may introduce bias when comparing the extracted local graph structures to determine ”frequency.” However, nodes in the available action space are always unselected, meaning they share the same state value in $\bm{x}_{t}$. Consequently, this concatenation does not significantly impact the ranking of each action’s probability or value. Also, for smaller values of $k$, only a few nodes are selected, which does not lead to much bias when comparing local graph structures, and the bias increases as $k$ becomes larger. Therefore, there is no significant performance gap between two approaches for smaller values of $k$, but RLMiner with state-aware message passing performed better for larger value of $k$.